# Some Super-approximation Rates of ReLU Neural Networks for Korobov Functions

### v1

@@ Abstract
This paper examines the $L_{p}$ and $W^{1}_{p}$ norm approximation errors of ReLU neural networks for Korobov functions. In terms of network width and depth, we derive nearly optimal super-approximation error bounds of order $2m$ in the $L_{p}$ norm and order $22m-$ 2 in the $W^{1}_{p}$ norm, for target functions with $L_{p}$ mixed derivative of order $m$ in each direction. The analysis leverages sparse grid finite elements and the bit extraction technique. Our results improve upon classical lowest order $L_{\infty}$ and $H^{1}$ norm error bounds and demonstrate that the expressivity of neural networks is largely unaffected by the curse of dimensionality.
@@
[…]
[inst1]addressline=School of Mathematical Sciences, city=Zhejiang University, postcode=310058, state=Hangzhou, country=China
@@ 1 Introduction
In recent years, Deep Neural Networks (DNNs) have emerged as a crucial technology in the fields of machine learning and artificial intelligence and achieved tremendous success. The remarkable success of DNNs has inspired numerous practical applications in computer science and engineering, including image recognition, natural language processing, autonomous driving, scientific computing, etc. Towards an explainable framework of machine learning, understanding the approximation power of DNNs is a fundamental question and remains an active research area. In particular, one challenge in this area is characterizing the approximation error bound for certain function class in terms of network depth and width, see, e.g., Yarotsky (2018); Yarotsky and Zhevnerchuk (2020); Shen et al. (2019); Lu et al. (2021); Shen et al. (2020, 2021a); DeVore et al. (2021); Zhang et al. (2022); He et al. (2022); Zhang et al. (2024); Daubechies et al. (2022); Siegel and Xu (2022); Siegel (2023); Yang et al. (2023a); Yang (2024); He (2024); Li and Siegel (2024).
Recently, the bit extraction technique has gained renewed attention for verifying if certain functions exhibit super-approximation rates using well-designed network architectures. This technique was invented by Bartlett et al. (1998); Anthony and Bartlett (1999) to compute VC-dimension bounds of DNNs, as discussed in Blumer et al. (1989); Harvey et al. (2017); Bartlett et al. (2019). In Yarotsky (2018); Shen et al. (2020); Lu et al. (2021); Siegel (2023); Liu et al. (2024a); Yang (2024), the bit extraction technique was further used to produce nearly optimal super-approximation error bounds of DNNs for continuous functions and Sobolev functions, while the order of approximation rate deteriorates as the dimension grows.
To partially cure the curse of dimensionality, Korobov regularity has been extensively applied in approximation error analysis of neural networks, with a focus on shallow networks (Blanchard and Bennouna (2022); Liu et al. (2024b)), fully connected deep neural networks (Montanelli and Du (2019); Yang and Lu (2024)), convolutional deep neural networks (Mao and Zhou (2022); Li and Zhang (2025); Fang et al. (2025)), and quantum circuits (Aftab and Yang (2024)). Analysis in the aforementioned works relies on interpolation error estimates on sparse grids (Bungartz and Dornseifer (1998); Bungartz and Griebel (2004)). We shall follow this line and derive novel near-optimal approximation error bounds of arbitrary order for Korobov functions.
@@ 1.1 Main Results
Prior to this work, Yang and Lu (2024) has derived approximation error bounds for Korobov functions (see Definition 1) in $X^{2}_{\infty}(\Omega)$ using DNNs with a Rectified Linear Unit (ReLU) activation $\sigma(x):=\max(x,0)$, measured in the $L_{p}(\Omega)$ norm and the $W^{1}_{2}(\Omega)$ Sobolev norm.
Definition 1 (Korobov space).
@@
[…]
@@
$|f|_{m,p}:=\left\|\frac{\partial^{md}f}{\partial x_{1}^{m}\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu}\partial x_{d}^{m}}\right\|_{L_{p}(\Omega)}.$
A natural question is whether increasing the smoothness of the target function leads to improved approximation rates. In Theorem 1, we establish higher order super-approximation $L_{p}$ error bounds of ReLU DNNs for Korobov functions in $X^{m}_{p}(\Omega)$.
Theorem 1.
Given any $f\in X^{m}_{p}(\Omega)$ with $1\leq p<\infty$ and $m\geq 2$, for any $W,L\in\mathbb{N}_{+}$, there is a function $\phi$ implemented by a ReLU DNN with width $C_{1}W(\log(8W))^{d+1}$ and depth $C_{2}L(\log(8L))^{d+1}$ such that
$\|f-\phi\|_{L_{p}(\Omega)}\leq C_{3}|f|_{m,p}W^{-2m}L^{-2m}(\log(2WL))^{3d},$
where $C_{1}=112d(2d)^{d}$, $C_{2}=320d^{2}$, and $C_{3}=(2d)^{3d}$ if $m=2$; $C_{1}=70md^{d+1}$, $C_{2}=90m^{2}d^{2}2^{d}$, and $C_{3}$ is a constant depending only on $m$ and $d$ if $m\geq 3$.
The authors of Yang and Lu (2024) conjectured that the super-approximation error bound of ReLU DNN for target functions in $X^{2}_{p}(\Omega)$ is $O(W^{-4+\frac{1}{p}}L^{-4+\frac{1}{p}})$ (up to logarithmic factors). The lowest order version of Theorem 1 disproves this conjecture and illustrates that the ReLU DNN achieves the nearly optimal approximation error bound regardless of the integrability index $p$.
@@
Theorem 2.
Given any $f\in X^{m}_{p}(\Omega)$ with $1\leq p<\infty$ and $m\geq 2$, for any $W,L\in\mathbb{N}_{+}$, there is a function $\phi$ implemented by a ReLU DNN with width $C_{4}W(\log(8W))^{d+1}$ and depth $C_{5}L(\log(8L))^{d+1}$ such that
$\|f-\phi\|_{W^{1}_{p}(\Omega)}\leq C_{6}|f|_{m,p}W^{-2(m-1)}L^{-2(m-1)}(\log(2WL))^{3d},$
where $C_{4}=89d(2d)^{2d}m^{2}$, $C_{5}=97d^{2}m^{2}$, and $C_{6}$ is a constant depending only on $m,p$ and $d$.
We can further establish that the error bounds of ReLU DNNs in Theorems 1 and 2 are nearly optimal: For any $\delta,C>0$, there exist $f\in X^{m}_{p}(\Omega)$ with $|f|_{m,p}\leq 1$ and $W,L\in\mathbb{N}_{+}$ such that
@@
This result can be proved using (Siegel, 2023, Theorem 22) and the VC-dimension lower bound of ReLU DNNs (Bartlett et al. (2019); Yang et al. (2023b)).
The rest of this paper is organized as follows. In Section 2, we provide preliminaries on sparse grids and several lemmas required to prove our main results. In Section 3 and 4, we prove Theorem 1 and 2, respectively. Section 5 is devoted to concluding remarks.
@@ 1. $\bullet$
    Let $\mathbb{N}_{+}$ denote the set of all positive integers, i.e., +=$\mathbb{N}_{+}=\{1,2,\ldots\}$ and let :=+$\mathbb{N}:=\mathbb{N}_{+}\cup\{0\}$.
@@ 2. $\bullet$
    A DNN function $\phi:\mathbb{R}^{d}\longrightarrow\mathbb{R}$ can be realized through the following function compositions
    $\phi=\mathcal{A}_{L}\circ\sigma\circ\mathcal{A}_{L-1}\circ\sigma\circ\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu}\sigma\circ\mathcal{A}_{1}\circ\sigma\circ\mathcal{A}_{0},$
    where $\mathcal{A}_{i}(\bm{x}_{i}):=\bm{W}_{i}\bm{x}_{i}+\bm{b}_{i}$, with $\bm{W}_{i}\in\mathbb{R}^{N_{i+1}\times N_{i}}$ and $\bm{b}_{i}\in\mathbb{R}^{N_{i+1}}$ for all $\bm{x}_{i}\in\mathbb{R}^{N_{i}}$. The activation function $\sigma$ acts on vectors in a component-wise way. In addition, we set $N_{0}=d,N_{L+1}=1$. And $W:=\max_{1\leq i\leq L}N_{i}$ and $L$ refer to the width and depth of the DNN function $\phi$, respectively.
@@ 4. $\bullet$
    Let $\mathbf{1}_{E}$ be the characteristic function on a set $E$, i.e.,
    $\mathbf{1}_{E}=\begin{cases}1,&x\in E,\\ 0,&x\notin E.\end{cases}$
@@ 7. $\bullet$
    For any multi-index $\bm{\alpha}=(\alpha_{1},\alpha_{2},\ldots,\alpha_{d})\in\mathbb{N}_{+}^{d}$, let
    $D^{\bm{\alpha}}:=\frac{\partial^{|\bm{\alpha}|}}{\partial x_{1}^{\alpha_{1}}\partial x_{2}^{\alpha_{2}}\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu}\partial x_{d}^{\alpha_{d}}}$
    and $|\bm{\alpha}|:=\alpha_{1}+\alpha_{2}+\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu}+\alpha_{d}$.

### v2

@@ Abstract
This paper examines the $L_{p}$ and $W^{1}_{p}$ norm approximation errors of ReLU neural networks for Korobov functions. In terms of network width and depth, we derive nearly optimal super-approximation error bounds of order $2m$ in the $L_{p}$ norm and order 2m22m-2 in the $W^{1}_{p}$ norm, for target functions with $L_{p}$ mixed derivative of order $m$ in each direction. The analysis leverages sparse grid finite elements and the bit extraction technique. Our results improve upon classical lowest order $L_{\infty}$ and $H^{1}$ norm error bounds and demonstrate that the expressivity of neural networks is largely unaffected by the curse of dimensionality.
@@
{AMS}
68T07, 65D40, 41A25
@@ 1 Introduction
In recent years, Deep Neural Networks (DNNs) have emerged as a crucial technology in the fields of machine learning and artificial intelligence and achieved tremendous success. The remarkable success of DNNs has inspired numerous practical applications in computer science and engineering, including image recognition, natural language processing, autonomous driving, scientific computing, etc. Towards an explainable framework of machine learning, understanding the approximation power of DNNs is a fundamental question and remains an active research area. In particular, one challenge in this area is characterizing the approximation error bounds for a certain function class in terms of network depth and width, see, e.g., [42, 41, 30, 26, 31, 32, 11, 44, 14, 43, 10, 34, 35, 37, 40, 15, 21].
Recently, the bit extraction technique has gained renewed attention for verifying if certain functions exhibit super-approximation rates using well-designed network architectures. This technique was invented by [4, 2] to compute VC-dimension bounds of DNNs, as discussed in [6, 13, 3]. In [42, 31, 26, 35, 24, 40], the bit extraction technique was further used to produce nearly optimal super-approximation error bounds of DNNs for continuous functions and Sobolev functions, while the order of approximation rate deteriorates as the dimension grows.
To partially cure the curse of dimensionality, Korobov regularity has been extensively applied in approximation error analysis of neural networks, with a focus on shallow networks ([5, 25]), fully connected deep neural networks ([28, 36]), convolutional deep neural networks ([27, 22, 12]), and quantum circuits ([1]). Analysis in the aforementioned works relies on interpolation error estimates on sparse grids ([8, 9]). We shall follow this line and derive novel near-optimal approximation error bounds of arbitrary order for Korobov functions.
@@ 1.1 Main Results
Prior to this work, [36] has derived approximation error bounds for Korobov functions (see Definition 1.1) in $X^{2}_{\infty}(\Omega)$ using DNNs with a Rectified Linear Unit (ReLU) activation $\sigma(x):=\max(x,0)$, measured in the $L_{p}(\Omega)$ norm and the $W^{1}_{2}(\Omega)$ Sobolev norm.
Definition 1.1 (Korobov space).
@@
$X^{m}_{p}(\Omega):=$ $\big\{f\in L_{p}(\Omega):D^{\bm{\alpha}}f\in L_{p}(\Omega)\text{ for each multi-index}$
@@
$|f|_{m,p}:=\left\|\frac{\partial^{md}f}{\partial x_{1}^{m}\cdots\partial x_{d}^{m}}\right\|_{L_{p}(\Omega)}.$
A natural question is whether increasing the smoothness of the target function leads to improved approximation rates. In Theorem 1.2, we establish higher order super-approximation $L_{p}$ error bounds of ReLU DNNs for Korobov functions in $X^{m}_{p}(\Omega)$.
Theorem 1.2.
Given any $f\in X^{m}_{p}(\Omega)$ with $1\leq p<\infty$ and $m\geq 2$, for any $W,L\in\mathbb{N}_{+}$, there is a function $\phi$ implemented by a ReLU DNN with width $W$ and depth $L$ such that
$\|f-\phi\|_{L_{p}(\Omega)}\leq\widetilde{C}_{1}|f|_{m,p}W^{-2m}L^{-2m}(\log(8W))^{2(m+2)(d+1)}(\log(8L))^{2(m+2)(d+1)},$
where $\widetilde{C}_{1}=(C_{1}C_{2})^{4}C_{3}$, $C_{1}=112d(2d)^{d}$, $C_{2}=320d^{2}$, and $C_{3}=(2d)^{3d}$ if $m=2$; $\widetilde{C}_{1}$ is a constant depending only on $m$ and $d$ if $m\geq 3$.
Sketch of proof: For any $f\in X^{m}_{p}(\Omega)$, the proof proceeds in three main steps. First, we utilize the sparse grid interpolation $\Pi_{n}^{m}f(\bm{x})=\sum_{|\bm{l}|_{1}\leq n+d-1}\sum_{\bm{i}\in\mathcal{I}_{\bm{l}}}v_{\bm{l},\bm{i}}^{m}\phi_{\bm{l},\bm{i}}^{m}(\bm{x})$, which is known to approximate $f$ with the desired error bounds. Second, for each level $\bm{l}$, we construct a ReLU DNN $\phi_{\bm{l}}$ to approximate the sub-sum $\sum_{\bm{i}\in\mathcal{I}_{\bm{l}}}v_{\bm{l},\bm{i}}^{m}\phi_{\bm{l},\bm{i}}^{m}$ within a subdomain $\Omega_{\varepsilon}\subset\Omega$. Specifically, since the basis functions $\phi_{\bm{l},\bm{i}}^{m}$ have disjoint supports for a fixed level $\bm{l}$, the sub-sum simplifies locally to a single term $v_{\bm{l},\bm{i}}^{m}\phi_{\bm{l},\bm{i}}^{m}$. Each such term is approximated by a network that integrates a product sub-network, a sub-network representing the basis function, and one approximating the coefficients. Finally, we aggregate these networks $\phi_{\bm{l}}$ for all $|\bm{l}|_{1}\leq n+d-1$ to form the ultimate neural network. For $p<\infty$, the approximation error on $\Omega_{\varepsilon}$ can be extended to the entire domain $\Omega$ by choosing a sufficiently small $\varepsilon$, thereby preserving the desired convergence rates.
The authors of [36] conjectured that the super-approximation error bound of ReLU DNN for target functions in $X^{2}_{p}(\Omega)$ is $O(W^{-4+\frac{1}{p}}L^{-4+\frac{1}{p}})$ (up to logarithmic factors). The lowest order version of Theorem 1.2 disproves this conjecture and illustrates that the ReLU DNN achieves the nearly optimal approximation error bound regardless of the integrability index $p$.
@@
Theorem 1.3.
Given any $f\in X^{m}_{p}(\Omega)$ with $1\leq p<\infty$ and $m\geq 2$, for any $W,L\in\mathbb{N}_{+}$, there is a function $\phi$ implemented by a ReLU DNN with width $W$ and depth $L$ such that
$\|f-\phi\|_{W^{1}_{p}(\Omega)}\leq\widetilde{C}_{2}|f|_{m,p}W^{-2(m-1)}L^{-2(m-1)}(\log(8W))^{2(m+1)(d+1)}(\log(8L))^{2(m+1)(d+1)},$
where $\widetilde{C}_{2}$ is a constant depending only on $m,p$ and $d$.
Sketch of proof: For any $f\in X^{m}_{p}(\Omega)$, we adopt a distinct methodology for the $W^{1}_{p}$ norm analysis. This approach bypasses the need for the domain-extension process (from $\Omega_{\varepsilon}$ to $\Omega$) typically required in the $L_{p}$ case for $p<\infty$. First, we utilize the sparse grid interpolation $\Pi_{n}^{m}f(\bm{x})=\sum_{|\bm{l}|_{1}\leq n+d-1}\sum_{\bm{i}\in\mathcal{I}_{\bm{l}}}v_{\bm{l},\bm{i}}^{m}\phi_{\bm{l},\bm{i}}^{m}(\bm{x})$, which is known to approximate $f$ with the desired error bounds in the $W^{1}_{p}$ norm. Second, we use a partition of unity $\{g_{\bm{k}}\}$ satisfying $\sum_{\bm{k}}g_{\bm{k}}\equiv 1$ on $\Omega=\cup\Omega_{\bm{k}}$, where each $g_{\bm{k}}$ is efficiently approximated on $\Omega_{\bm{k}}$ by a ReLU DNN $\phi_{\bm{k}}$. Within each subdomain $\Omega_{\bm{k}}$, we construct a ReLU DNN $\psi_{\bm{k}}^{\bm{l}}$ to approximate the sub-sum $\sum_{\bm{i}\in\mathcal{I}_{\bm{l}}}v_{\bm{l},\bm{i}}^{m}\phi_{\bm{l},\bm{i}}^{m}$ in both $W^{1}_{p}$ and $L_{p}$ norms, following a procedure analogous to the $L_{p}$ case. Next, we aggregate these sub-networks $\psi_{\bm{k}}^{\bm{l}}$ over all levels $|\bm{l}|_{1}\leq n+d-1$ to form a localized network $\psi_{\bm{k}}$ that approximates $f$ on $\Omega_{\bm{k}}$ by incorporating the interpolation error. Finally, the global approximation error on $\Omega$ is decomposed across the subdomains $\Omega_{\bm{k}}$ via the inequality $\|f-\phi\|\leq\sum_{\bm{k}}\|fg_{\bm{k}}-\widetilde{\phi}(\phi_{\bm{k}},\psi_{\bm{k}})\|$, where $\widetilde{\phi}$ denotes the product network derived from Lemma 2.6.
The approximation errors measured in the $L_{p}$ and $W^{1}_{p}$ norms demonstrate that ReLU DNNs can effectively double the approximation rate of conventional methods—such as continuous function approximators (cf. [36, 11, 35, 40])—with respect to the network depth $L$ and width $N$. This phenomenon is thus referred to as the super-approximation of ReLU DNNs. We can further establish that the error bounds of ReLU DNNs in Theorems 1.2 and 1.3 are nearly optimal: For any $\delta,C>0$, there exist $f\in X^{m}_{p}(\Omega)$ with $|f|_{m,p}\leq 1$ and $W,L\in\mathbb{N}_{+}$ such that
@@
This result can be proved using [35, Theorem 22] and the VC-dimension bound of ReLU DNNs ([3, 38]).
Remark 1.4.
By the definition of Korobov spaces, we observe the inclusion $W^{md}_{p}(\Omega)\cap\{f:f|_{\partial\Omega}=0\}\subset X^{m}_{p}(\Omega)$. This inclusion is strict, i.e., there exist functions $f\in X^{2}_{p}(\Omega)$ such that $f\notin W^{2d}_{p}(\Omega)$ for any $1\leq p\leq\infty$ (cf. [27]). Consequently, existing $L_{p}$ norm error bound $O((WL)^{-2s/d})$ of DNN approximation for $W^{s}_{p}(\Omega)$ functions (cf. [35, 39]) with $s=md$ cannot be directly applied to Korobov functions to achieve the error bound $O((WL)^{-2m})$ (ignoring logarithmic factors). The mixed regularity of Korobov functions necessitates a non-trivial DNN approximation theory, which mitigates the curse of dimensionality by leveraging sparse grid approximation.
1.2 Other Network Architectures
As indicated in Corollary 2 in [38], for the approximation of Sobolev functions $f\in W^{m}_{\infty}(\Omega)$ in the $W^{n}_{\infty}$ norm with $n\geq 2$ using ReLU-$\text{ReLU}^{2}$ DNN (where the activation function applied to each neuron can be chosen as ReLU or its square) with width $O(W\log W)$ and depth $O(L\log L)$, the superconvergence rates $O((WL)^{-2(m-n)/d})$ can be achieved by employing a smoother unit decomposition similar to $g_{\bm{m}}$, i.e., utilizing piecewise polynomials of degree $n$. This methodology can actually be used to obtain the approximation of higher order Korobov functions using ReLU-$\text{ReLU}^{2}$ DNNs in the norm of $W^{n}_{p}$ with $n\geq 2$, with the help of the error estimates of sparse grid interpolation in this norm. This error can actually be computed using the inverse estimates akin to the process of calculating the $W^{1}_{p}$ error.
Recently, Floor-ReLU networks (i.e., the activation function applied to each neuron can be chosen as ReLU or Floor function) with super approximation power have been introduced in [32] to efficiently approximate Hölder functions on $\Omega$. By observing that the proof process in [32, Proposition 3.2] have provided that for $K=W^{L}$ and $\xi_{i}\in[0,1]$ for $i=1,2,\ldots,K$, there exists a function $\psi\in\mathcal{NN}(2W^{2}+5W,L(7L-1)+2)$ such that $|\psi(i)-\xi_{i}|\leq 2^{-WL}$, we can provide the results for approximating Korobov functions on $\Omega$ by the Floor-ReLU DNN. Similar to ReLU DNN approximation in the norm of $W^{1}_{p}$, establishing a series of results about approximating polynomial by Floor-ReLU DNN in the $W^{1}_{p}$ norm will lead to the approximation results for higher order Korobov functions using Floor-ReLU DNNs, which will be left for future research.
Recently, the authors of [24] have shown that ResNet ([16, 23, 29, 20]) with constant width and depth $O(L)$ can approximate Lipschitz continuous function $f\in C([0,1]^{d})$ with the optimal approximation rates $O(\omega_{f}(L^{-2/d}))$, where $\omega_{f}$ is the modulus of continuity of the function $f$. Additionally, the versions of ResNet similar to Lemma 2.1, Lemma 2.3, and Lemma 2.7 for $L_{\infty}$ norm are presented. By combining these results with the representation of basis functions by ReLU networks as demonstrated in the proof of Theorem 1.2, one can derive the approximation rates for Korobov spaces using ResNet through analogous methodologies. For brevity, we will not delve into the proof here. Moreover, we can obtain approximation results of ResNet for Korobov functions in the $W^{1}_{p}$ norm, resembling those outlined in Theorem 1.3, by formulating the $W^{1}_{p}$ ResNet counterparts of Lemma 2.1, 2.3, and 2.7.
The rest of this paper is organized as follows. In Section 2, we provide preliminaries on sparse grids and several lemmas required to prove our main results. In Section 3 and 4, we prove Theorem 1.2 and 1.3, respectively. Section 5 is devoted to concluding remarks.
@@ 1. $\bullet$
    Let $\mathbb{N}_{+}$ denote the set of all positive integers, i.e., $\mathbb{N}_{+}=\{1,2,\ldots\}$ and let $\mathbb{N}:=\mathbb{N}_{+}\cup\{0\}$.
@@ 2. $\bullet$
    For any $x\in\mathbb{R}$, let $\lceil x\rceil:=\min\{n:n>x,n\in\mathbb{Z}\}$.
  3. $\bullet$
    A DNN function $\phi:\mathbb{R}^{d}\longrightarrow\mathbb{R}$ can be realized through the following function compositions ϕ=A_L∘σA_L-1∘σ∘⋯σA_1∘σA_0, where $\mathcal{A}_{i}(\bm{x}_{i}):=\bm{W}_{i}\bm{x}_{i}+\bm{b}_{i}$, with $\bm{W}_{i}\in\mathbb{R}^{N_{i+1}\times N_{i}}$ and $\bm{b}_{i}\in\mathbb{R}^{N_{i+1}}$ for all $\bm{x}_{i}\in\mathbb{R}^{N_{i}}$. The activation function $\sigma$ acts on vectors in a component-wise way. In addition, we set $N_{0}=d,N_{L+1}=1$. And $W:=\max_{1\leq i\leq L}N_{i}$ and $L$ refer to the width and depth of the DNN function $\phi$, respectively.
@@ 5. $\bullet$
    Let $\mathbf{1}_{E}$ be the characteristic function on a set $E$, i.e., 1_E= {1, x ∈E,0, x∉E.
@@ 8. $\bullet$
    For any multi-index $\bm{\alpha}=(\alpha_{1},\alpha_{2},\ldots,\alpha_{d})\in\mathbb{N}_{+}^{d}$, let D^α:=∂α∂x1α1∂x2α2⋯∂xdαd and $|\bm{\alpha}|:=\alpha_{1}+\alpha_{2}+\cdots+\alpha_{d}$.